OpenStudy (starryknight):

please help....explain in steps Solve by factoring. 4x2 - 8x + 3 = 0 A: x = 1.5 or x = 0.5 B: x = -1.5 or x = -0.5 C: x = 1.5 or x = -0.5 D: x = -2 or x = 1.5

OpenStudy (anonymous):

4x^2 - 8x + 3 = 0 (2x - 1)(2x - 3) = 0 2x - 1 = 0 x = 1/2 or .5 Do you see how I got this?

OpenStudy (starryknight):

how did you get (2x-1)(2x-3)

OpenStudy (anonymous):

\[4x ^{2}-8x+3=0\]

OpenStudy (anonymous):

Thats the equation right?

OpenStudy (starryknight):

yes

OpenStudy (anonymous):

Okay, you have to factor. What are two numbers you can get to make \[4x ^{2}?\]

OpenStudy (starryknight):

i honestly dont now :(

OpenStudy (anonymous):

Okay, that's fine. 2x times 2x equals what?

OpenStudy (starryknight):

4x^2

OpenStudy (anonymous):

Great job! Now we need to find what multiply to 3, but add up to -8. (2x^2-?)(2x^2-?) We know that the two numbers need to be negative because of that -8.

OpenStudy (anonymous):

 these numbers would work because we factor out the numbers.